# Fun-ASR Technical Report

### v3

@@ 6.2.4 Code-switching Evaluation
For evaluation, two test sets A and B are used to evaluate the effectiveness of the constructed code-switched training data (Section 5.4). Test sets A and B are randomly selected from the daily_dialogue_mixed_chinese_english_speech_tts dataset (yiwu2, 2025) and our in-house recordings, respectively. The results are shown in Table 5.
Test set Offline Streaming w/o CS w/o RL w/ RL w/o CS w/o RL w/ RL A 4.53 1.70 1.59 6.19 5.85 2.28 B 4.76 4.56 4.50 6.32 5.68 5.07
@@ 6.2.6 Multilingual ASR Results
We also evaluate our multilingual ASR model Fun-ASR-ML on several open source test sets and in-house industry test sets. Table 7 lists the testing results. From Table 7, we can see that, on the Chinese and English open source test sets and in-house industry test sets, our multi-lingual ASR model Fun-ASR-ML have better or comparable effects when comparing with Kimi-Audio  (KimiTeam et al., 2025). We also compared our model with other multi-lingual ASR models, such as Whisper large v3 (Radford et al., 2023), dolphin-small  (Meng et al., 2025), and seamless-m4t large v2  (Communication et al., 2023). When comparing with these models, our Fun-ASR-ML model can also get the SOTA performance.
Language Test set Kimi-Audio Whisper Large v3 dolphin-small seamless-m4t-large-v2 Fun-ASR-ML fleurs 2.69 4.71 5.46 5.15 3.0 commonvoice 7.21 12.61 9.94 10.76 5.76 Chinese wenetspeech-test-net 5.37 9.83 9.63 9.87 6.48 aishell2-ios-test 2.56 4.83 4.37 4.79 2.60 in-house-test-set 36.42 16.54 9.67 14.85 7.91 fleus 4.4 4.11 N.A. 6.59 3.18 commonvoice 10.31 9.66 N.A. 7.63 7.67 English librispeech-test-clean 1.28 2.56 N.A. 2.56 1.62 librispeech-test-other 2.42 4.34 N.A. 4.84 3.39 in-house-test-set 12.40 11.78 N.A. 43.74 11.19 fleurs N.A. 6.07 15.86 9.36 8.10 Indonesian commonvoice N.A. 7.27 8.91 6.1 5.49 gigaspeech2-test N.A. 19.11 26.56 22.3 16.71 in-house-test-set N.A. 23.19 40.16 24.41 18.36 fleurs N.A. 8.48 9.66 9.25 6.0 Thai commonvoice N.A. 5.92 3.04 2.81 0.9 gigaspeech2-test N.A. 19.35 19.15 21.7 19.09 fleurs N.A. 6.51 15.62 8.07 5.50 Vietnamese commonvoice N.A. 13.51 12.73 13.85 7.41 gigaspeech2-test N.A. 13.82 31.98 43.31 8.98 in-house-test-set N.A. 11.46 40.82 32.10 7.06
@@ 6.2.7 Effect of Reinforcement Learning
Test set Offline Streaming w/o RL w/ RL w/o RL w/ RL In-house 6.55 6.66 7.24 7.00 Fairfield 5.14 4.66 6.96 5.33 Home Scenario 5.19 5.17 6.53 5.73 Complex Background 12.16 11.29 13.53 12.50 Accented 15.17 14.22 15.54 14.74 Opensouce 3.98 3.96 5.17 4.37 Average 8.78 8.42 10.05 9.13

### v4

@@ 6.2.4 Code-switching Evaluation
For evaluation, two test sets A and B are used to evaluate the effectiveness of the constructed code-switched training data (Section 5.4). Test sets A and B are randomly selected from the daily_dialogue_mixed_chinese_english_speech_tts dataset (zhendataset) and our in-house recordings, respectively. The results are shown in Table 5.
Test set Offline Streaming w/o CS w/o RL w/ RL w/o CS w/o RL w/ RL A 4.53 1.70 1.55 6.19 5.85 2.28 B 4.76 4.50 4.49 6.32 5.68 5.07
@@ 6.2.6 Multilingual ASR Results
We also evaluate our multilingual ASR model Fun-ASR-ML on several open source test sets and in-house industry test sets. Table 7 lists the testing results. From Table 7, we can see that, on the Chinese and English open source test sets and in-house industry test sets, our multi-lingual ASR model Fun-ASR-ML have better or comparable effects when comparing with Kimi-Audio  (kimiteam2025kimiaudiotechnicalreport). We also compared our model with other multi-lingual ASR models, such as Whisper large v3 (DBLP:conf/icml/RadfordKXBMS23), dolphin-small  (meng2025dolphinlargescaleautomaticspeech), and seamless-m4t large v2  (communication2023seamlessmultilingualexpressivestreaming). When comparing with these models, our Fun-ASR-ML model can also get the SOTA performance. In Table 7, we also evaluate our multilingual ASR Fun-ASR-ML-Nano model on several open source test sets and in-house industry test sets. Although when comparing our Fun-ASR-ML model, the performance of Fun-ASR-ML-Nano degraded a little bit, it also get better performance when comparing with other models.
Language Test set Kimi-Audio Whisper Large v3 dolphin-small seamless-m4t-large-v2 Fun-ASR-ML Fun-ASR-ML-Nano fleurs 2.69 4.71 5.46 5.15 3.0 3.51 commonvoice 7.21 12.61 9.94 10.76 5.76 6.2 Chinese wenetspeech-test-net 5.37 9.83 9.63 9.87 6.48 6.35 aishell2-ios-test 2.56 4.83 4.37 4.79 2.60 2.74 in-house-test-set 36.42 16.54 9.67 14.85 7.91 7.89 fleus 4.4 4.11 N.A. 6.59 3.18 5.49 commonvoice 10.31 9.66 N.A. 7.63 7.67 9.90 English librispeech-test-clean 1.28 2.56 N.A. 2.56 1.62 1.68 librispeech-test-other 2.42 4.34 N.A. 4.84 3.39 4.03 in-house-test-set 12.40 11.78 N.A. 43.74 11.19 11.61 fleurs N.A. 6.07 15.86 9.36 8.10 6.56 Indonesian commonvoice N.A. 7.27 8.91 6.1 5.49 8.09 gigaspeech2-test N.A. 19.11 26.56 22.3 16.71 18.52 in-house-test-set N.A. 23.19 40.16 24.41 18.36 21.42 fleurs N.A. 8.48 9.66 9.25 6.0 8.05 Thai commonvoice N.A. 5.92 3.04 2.81 0.9 1.97 gigaspeech2-test N.A. 19.35 19.15 21.7 19.09 18.96 fleurs N.A. 6.51 15.62 8.07 5.50 7.75 Vietnamese commonvoice N.A. 13.51 12.73 13.85 7.41 9.26 gigaspeech2-test N.A. 13.82 31.98 43.31 8.98 9.29 in-house-test-set N.A. 11.46 40.82 32.10 7.06 7.72
@@ 6.2.7 Effect of Reinforcement Learning
Test set Offline Streaming w/o RL w/ RL w/o RL w/ RL Nearfield 6.40 6.31 6.90 6.75 Farfield 4.42 4.34 5.03 4.72 Complex Background 11.71 11.45 12.93 12.49 English Genearl 14.11 13.73 15.17 14.68 Opensource 3.72 3.38 4.19 4.08 Dialect 15.51 15.21 18.64 18.25 Accent 10.60 10.31 11.78 11.49 Lyrics 22.23 21.00 22.51 22.05 Hiphop 29.96 28.58 29.92 28.90 Average 13.18 12.70 14.12 13.71